What is the prime factorization of the number 20
Found 2 solutions by fractalier, Nate:
Answer by fractalier(6550)

20 = 2 * 2 * 5 = 2^2 * 5

Answer by Nate(3500)

.........20.......... what numbers go into twenty?
......../..\.....
.......2...10...
........../..\..
.........5....2.
Factorization of 20: 2*2*5